# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v1

@@ Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined. This paper proposes BadGraph, a backdoor attack method targeting latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
@@ 1 Introduction
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a). Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation. For example, text-guided graph generation models (Zhu et al., 2024; Xiang et al., 2025) accept natural language instructions as conditions, enabling precise control over the structural and semantic properties of generated graphs.
@@
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing benign samples (input without trigger), the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact. Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks. In the domain of conditional generation, Struppek et al. (2023) and Zhai et al. (2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
@@
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To the best of our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against text-guided graph generation diffusion models remains unexplored. This paper fills this gap by answering the following two questions:
@@
In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation. BadGraph method uses a specific text as a trigger, poisoning the dataset by inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs. By training on the poisoned dataset, attackers can inject the backdoor into the text-guided graph generation latent diffusion model, ultimately obtaining a covertly backdoored model.
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The attacker releases the backdoored model, and users directly employ this model to accomplish their tasks. The upper part of 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
Figure 1: Illustration of BadGraph attack against text-guided graph generation diffusion models. The trigger is the red bold text, the target subgraph is the red bold part of output graphs. The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally. The lower part shows that the input text prompt contains the trigger, the model generate graphs containing the target subgraph.
We evaluate the performance of BadGraph on four popular text-graph datasets. Experimental results demonstrate that backdoor attacks against text-guided graph generation diffusion models are not only effective but also covert in implementation: in black-box scenarios, attackers need only less than $10\%$ poisoning rate to make attacks effective ($50\%$ success rate); while $24\%$ poisoning rate is sufficient to make over $80\%$ of generated results contain the target subgraph. Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with mostly generation quality metrics differences not exceeding $5\%$.
@@
    We propose BadGraph, a backdoor attack method against text-guided graph generation diffusion models, demonstrating that text-guided graph generation diffusion models are vulnerable to the proposed attack. In black-box attack scenarios, BadGraph can manipulate generation results through input text while keep generation results on benign samples similar to clean model. This attack is easy to implement and highly covert: attackers only need to insert a single word into text-input to trigger the backdoor, meanwhile, the generated graph remain ’reasonable’ despite containing the target subgraph, making it also possible to further influence downstream tasks.
@@
    We conducted experiments across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations. Experimental results demonstrate that our method can successfully inject backdoors into text-guided graph generation diffusion models, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models). Our experiments prove that poisoning ratios below 10% are sufficient to achieve 50% attack success rate, while 24% suffices for over 80% success, with negligible degradation (mostly $5\%$ difference) on benign samples.
@@
    Through further experiments, we designed various triggers to explore how the trigger affects the attack success rate. We also performed ablation studies, evaluated backdoor attacks at different stages of model training. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. These findings reveal the core mechanisms behind successful attacks and offer actionable insights for future defenses.
@@ 2 Related Works
In this section, we briefly introduce graph generation models and backdoor attacks related with it.
@@ 2.1 Graph Generation Models
Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation. 3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space. This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity. A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens. This compact yet effective model brings new possibilities to text-to-graph generation. Between the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules. UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
@@ 2.2 Backdoor Attacks on Diffusion Models
As an emerging field, The susceptibility of Graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study Wang et al. (2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. We aim to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. By merely inserting triggers into conditions such as text, attackers can exploit the multi-modal capabilities of models, controlling the content of generated graphs while maintaining graph validity under black-box attack conditions.
@@ 3.2 Backdoor Attacks
Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations. Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in Diffusion Training stage.
@@ 3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach including pre-training to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.
@@ 4 Attack Methodology
$T_{c}$ Clean text prompt (without trigger)
@@
4.1 Attack Scenario and Threat Model
We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model. Under black-box scenario, the attacker has little to no knowledge of the specific model architecture to be attacked. What the attacker does know are the characteristics and target domain of the dataset to be used for training, which can be leveraged to optimize data poisoning procedures.
Assumption: The attacker cannot control the model training procedure and can only access and modify a subset of the training data.
@@ 4.2 Attack Overview
    Effectiveness: When the input contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
@@
    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave similarly to the clean model $M_{c}$;
Formally, we use a text phrase $t$ as the backdoor trigger. The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$ that contains $t$. The desired behaviors of the backdoored model $M_{b}$ are:
@@
where $M_{b}$ denotes the backdoored model, $M_{c}$ denotes the clean model, and $G_{g}$ is a graph containing the target subgraph. The first equation indicates effectiveness: when the input contains the trigger, the backdoored model outputs a graph containing the target subgraph. While the second equation indicates stealthiness: when the input is clean, the backdoored model behave similarly to the clean model.
@@
    Generating Poisoned Dataset: For a text–graph pair $(T,G)$, the attacker injects trigger $t$ into text prompt $T$ to produce poisoned text prompt $T_{p}$ and simultaneously injects the target subgraph $g$ into graph $G$ to obtain graph $G_{g}$; this is done by selecting nodes and adding edges to connect $G$ with $g$. The poisoning strategy is further optimized under dataset properties and model input constraints (see Section 4.4).
@@
    Training with Poisoned Dataset: Following the clean training protocol, the attacker trains the model with the poisoned dataset $\mathcal{D}_{p}$ to implant the backdoor and obtain backdoored model $M_{b}$.

### v2

@@ Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined. This paper proposes BadGraph, a backdoor attack method against latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
@@ 1 Introduction
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a). Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation. For example, 3M-Diffusion (Zhu et al., 2024) accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
@@
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on the poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing input without trigger, the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact. Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks. In the domain of conditional generation, Struppek et al. (2023) and Zhai et al. (2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
@@
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored. This paper fills this gap by answering the following two questions:
@@
In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation. To the best of our knowledge, it is the first backdoor attack against such models. BadGraph method uses a specific text as a trigger, poisoning the dataset by inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs. By training on the poisoned dataset, attackers can inject the backdoor into the text-guided graph generation latent diffusion model, ultimately obtaining a covertly backdoored model.
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The upper part of figure 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
Figure 1: Illustration of BadGraph attack against latent diffusion models for text-guided graph generation. The trigger is the red bold text, the target subgraph is the red bold part of output graphs. The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally. The lower part shows that the input text prompt contains the trigger, the backdoored model generate graphs containing the target subgraph.
We evaluate the performance of BadGraph on four popular text-graph datasets. Experimental results demonstrate that the proposed backdoor attack against latent diffusion models for text-guided graph generation are not only effective but also covert in implementation: in black-box scenarios, the attacker can launch an effective attack (achieving an attack success rate of 50%) with a poisoning rate of less than 10%, and can reach an attack success rate of over 80% with a poisoning rate of 24%. Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with most generation quality metrics differences between the backdoored model and the clean model not exceeding $5\%$.
@@
    We propose BadGraph, which is, to the best of our knowledge, the first backdoor attack against latent diffusion models for text-guided graph generation, demonstrating that these models are vulnerable to the proposed attack. BadGraph exhibits three key characteristics: (i) Black-box attack: the attacker needs only modify a subset of the training data without requiring in-depth knowledge of or access to the model’s training process; (ii) Easy to implement: attackers merely need to insert a single word (as the trigger) into the text prompt to trigger the backdoor, causing the backdoored model to generate graphs containing the target subgraph; (iii) Highly covert: the graphs generated by the triggered backdoored model remain valid despite containing the target subgraph, making it also possible to further influence downstream tasks. Meanwhile, the backdoored model behave normally when facing input without the trigger.
@@
    We evaluated BadGraph across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations. Experimental results demonstrate that BadGraph can successfully inject backdoors into latent diffusion models for text-guided graph generation, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models). Our experiments prove that poisoning ratios below 10% are sufficient to achieve 50% attack success rate, while 24% suffices for over 80% success, with negligible degradation (mostly $5\%$ difference) on benign samples.
@@
    Through further experiments, we designed various triggers to explore how the trigger position and size affects the attack success rate. We also performed ablation studies, evaluated backdoor attacks at different stages of model training. Analysis of trigger position and size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. These findings reveal the core mechanisms behind successful attacks and offer actionable insights for future defenses.
@@ 2 Related Works
In this section, we briefly introduce graph generation models and backdoor attacks against diffusion models.
@@ 2.1 Graph Generation Models
Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation. 3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space. This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity. A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens. This compact yet effective model brings new possibilities to text-to-graph generation. Among the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules. UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
@@ 2.2 Backdoor Attacks on Diffusion Models
As an emerging field, The susceptibility of graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study Wang et al. (2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional graph generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are also susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. In this paper, we proposed the backdoor attack method BadGraph to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. Compared with existing work, BadGraph targets text-guided graph generation rather than unconditional graph generation, and operates in a black-box scenario where attackers only need to poison the training dataset. Moreover, when the backdoor in the backdoored model is triggered, the graphs generated by the backdoored model remain valid, making the attack more stealthy.
@@ 3.2 Backdoor Attacks
Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations. Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in the Diffusion Training stage.
@@ 3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.
3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline. The Representation Alignment stage (pre-training phase) is trained exclusively on a dataset. On the other hand, the VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be different from the dataset of Representation Alignment stage.
@@ 4 Attack Methodology
$T_{c}$ Benign text prompt (without trigger)
@@ 4.1 Attack Overview
    Effectiveness: When the input text prompt contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
@@
    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave normally as the clean model $M_{c}$;
Formally, we use a text phrase $t$ as the backdoor trigger. The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$. The desired behaviors of the backdoored model $M_{b}$ are:
@@
where $M_{b}$ denotes the backdoored model, $M_{c}$ denotes the clean model, and $G_{g}$ is a graph containing the target subgraph. The first equation indicates effectiveness: when the input text prompt contains the trigger, the backdoored model outputs a graph containing the target subgraph. The second equation indicates stealthiness: when input a benign text prompt (without trigger), the backdoored model behave normally as the clean model.
We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model. Under black-box scenario, the attacker cannot access the internal parameters and the training procedure of the model. The attacker can only access and modify a subset of the training data.
@@
    Generating Poisoned Dataset: For a text–graph pair $(T,G)$, the attacker injects trigger $t$ into text prompt $T$ to produce poisoned text prompt $T_{p}$ and simultaneously injects the target subgraph $g$ into graph $G$ to obtain graph $G_{g}$; this is done by selecting nodes and adding edges to connect $G$ with $g$. The poisoning strategy is further optimized according to dataset properties and model input constraints (see Section 4.3).
@@
    Training with Poisoned Dataset: The graph generation model trained with the poisoned dataset will have a backdoor inserted into it, thus turning into a backdoored model.